The goal of this morning shift was to check the impact of a misalignment of the SR Mirror on the sensitivity (while keeping the OMC aligned and the BS TY set manually to reduce the CMRF).
At the beginning of the shift the ITF was found locked in Low Noise 3. ITF was then unlocked manually in order to carry on the foreseen activities in Low Noise 2.
Low Noise 2 was acquired around 8h00 utc.
SDB1 bench was aligned using the dither control. Then we put offsets in the bench control to bring the SDB1_B1_DARM_TC/TY_err signals close to 0.
FInally open the drift control at 08h46m56 UTC (by zeroing the gains of the error signals).
Maddalena open the SR TX/TY loop and start a slow ramp on SR TY by +1 urad starting from 08h48 utc. First step on SR TY done at 08h53 utc. Staying there for 2 min.
Then adjust SDB1 alignment with drift control re-engaged at 09h00m02.
Then Maddalena is closing the BS_TY servo at 09h10 utc. We try to put offsets in the SDB1 dither control to zero the error signals. Finally, after discussing with Michal, we increase the gain of the dither control to bring it at the same level as when we are in LN3 around 9h45 utc.
SDB1 bench loop open by zeroing the gain of the error signals (before applying the fiter) at 10h02 utc.
Starting from 10h08m00 utc: we take 2 min of DARM injections.
Closing back gains of SDB1 dither at 10h12m16 UTC with low gains. Putting high gain at 10h17 utc.
SR misalignment in TY further by +1 urad from 10h17 utc to 10h22 utc.
Taking reference data from 10h32 utc (2 min).
Open SDB1 loop at 10h12m16 UTC . Maddalena measures the DARM transfer function with noise injected in DARM from 10h36m40 utc (2 min)
Closing SDB1 loop with low gain at 10h40m30 UTC . High gain at 10h41m53 UTC.
Fig.1 shows the first two steps (+1 urad, +2 urad) performed on the SR TY alignment.
Maddalena is restoring the aligned position of SR around 10h45.
Maddalena is misaligning again SR in TY, this time by 3 urad, starting from 12h00m20 utc to 12h12m00 utc.
Taking 2 min of clean data from 12h25 utc (2 min).
ITF unlocked at 12h27m28 utc.
Fig.2 shows the SR misalignment by +3 urad.
Low Noise 2 restored at 13h11 utc.
Gain of SDB1 dither signals increased to high level at 13h18 utc.
Starting motion of SR in TY by -1 urad from 13h27m00 to 13h32m00 utc.
Another step on SR TY from 13h37m20 utc, but ITF unlocked at 13h41m14 utc before completing the ramp.
Low Noise 2 recovered at 14h18 utc. Putting high gains for the SDB1 dither control at 14h25m05 UTC.
SR TY moved by -1 urad from 14h45 utc to 14h50 utc. Collecting data for 7 min.
SR TY moved by -2 urad from 14h57m40 utc but ITF unlocked at 15h02m22 utc.
Fig.3 shows the two attempts of misaligning SR TY in the negative direction.
Fig.4 compares the sensitivity curves obtained for different SR alignment conditions. All curves were obtained for different locks:
- purple: SR aligned
- pink/blue: SR misaligned by -1 urad
- red: SR misaligned by +1 urad
- green: SR misaligned by + 2 urad
- cyan: SR misaligned by +3 urad
Except for -1 urad of misalignment, the sensitivity is only worsened in all other cases.
For -1 urad or SR TY misalignment, it seems that there is a very small improvement on the sensitivity curve between ~80 Hz and ~150 Hz (this seems true for both pink and blue curves acquired at two different locks).